OpenStudy (lukecrayonz):

Evaluate. Integral s^5(2+s^6)^8ds

OpenStudy (lukecrayonz):

\[\int\limits_{?}^{?}s^{5}(2+s^{6})^{8}\]

OpenStudy (turingtest):

what u substitutions have you tried?

myininaya (myininaya):

Not always, but usually a good algebraic substitution to start with is trying let u equal the inside of a nasty part ( if that makes sense)

OpenStudy (lukecrayonz):

None since I have never learned this haha, this is my first question!:-)

OpenStudy (johnweldon1993):

Oh, okay...well, do you see a function in this integral where the derivative of that function is also shown?

OpenStudy (lukecrayonz):

It's kinda like f(g(x))g'(x) if that makes sense lol

OpenStudy (johnweldon1993):

Yeah something like that lol...so here I'll kind of give you this one... A good place to start, is to try an take the innermost function and see if the derivative (or some variation of it) is also shown in the integral...so here... The innermost would be (2 + s^6) what is the derivative of that function?

OpenStudy (lukecrayonz):

6s^5

OpenStudy (lukecrayonz):

@myininaya @TuringTest what about for this one? \[\int\limits_{a}^{b}\frac{ lnx^{2} }{ 5x }dx\]

myininaya (myininaya):

what is the derivative of ln(x)?

OpenStudy (lukecrayonz):

1/x

myininaya (myininaya):

you should see 1/x in your problem

myininaya (myininaya):

also I'm guessing that square is on ln(x) and not just the x am i wrong or right?